# Just ASR + LLM? A Study on Speech Large Language Models’ Ability to Identify and Understand Speaker in Spoken Dialogue

### v1

@@ 6 Conclusion
In this work, we categorize speaker-related spoken dialogue QA (SQA) questions into Identity-Critical Questions (ICQ) and Context-Based Questions (CBQ). We hypothesize that current state-of-the-art speech large language models (SpeechLLMs) have limited capability in distinguishing speakers’ voice characteristics and thus cannot solve ICQs as well as CBQs. To support our hypothesis, we propose an automatic classification approach for analyzing ICQs and CBQs in large datasets, and conduct extensive experiments and analysis using Qwen-Audio and WavLLM on the Gaokao and ”What Do You Like?” datasets. Our results show that despite WavLLM and Qwen-Audio obtain high accuracy on speaker-related questions in Gaokao, their performance on the ICQ portions is significantly poorer. This highlights the need for SpeechLLMs with enhanced speaker understanding capability as well as more comprehensive training and evaluation data for spoken dialogue understanding.

### v2

@@ 6 Conclusion
In this work, we categorize speaker-related spoken dialogue QA (SQA) questions into Identity-Critical Questions (ICQ) and Context-Based Questions (CBQ). We hypothesize that current state-of-the-art speech large language models (SpeechLLMs) have limited capability in distinguishing speakers’ voice characteristics and thus cannot solve ICQs as well as CBQs. To support our hypothesis, we propose an automatic classification approach for identifying ICQs and CBQs in large datasets, and conduct experiments and analysis using Qwen-Audio, WavLLM and text-only models on the Gaokao and What Do You Like? datasets. Our results show that WavLLM and Qwen-Audio obtain much lower accuracy on the speaker-related ICQ subset of Gaokao. Further, in the controlled What Do You Like? experiments, WavLLM performed similarly or worse than its text-only equivalent, supporting our hypothesis.
For future work towards improving SpeechLLM capabilities, it may be useful to explore alternative training techniques, e.g. pre-training with same/different speaker prediction tasks or RLHF with tasks that require speaker information. In addition, we argue that it is important to create a new SQA training and evaluation corpus that explicitly requires speaker identification capability and explores more attributes than gender.